OpenStudy (anonymous):

A 7ft thick slice is cut off the top of a cube, resulting in a rectangular box that has volume 37 ft^3. find the side length of the original cube

jimthompson5910 (jim_thompson5910):

Let s = side length of cube

jimthompson5910 (jim_thompson5910):

the original cube has a volume of V = s^3

jimthompson5910 (jim_thompson5910):

The part you cut off that is 7 ft high has a volume of 7s^2, since you multiply height by area of the base

jimthompson5910 (jim_thompson5910):

So Volume of Remaining Piece = (Original Volume) - (Volume of Cut Off Piece) Volume of Remaining Piece = (s^3) - (7s^2) 37 = s^3 - 7s^2 0 = s^3 - 7s^2 - 37 s^3 - 7s^2 - 37 = 0 I'll let you finish

OpenStudy (anonymous):

im confused

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

how to get the side from what you have said

jimthompson5910 (jim_thompson5910):

well the best way to solve s^3 - 7s^2 - 37 = 0 is to use a graphing calculator to approximate the root

jimthompson5910 (jim_thompson5910):

you can't solve this exactly (not easily anyway)

jimthompson5910 (jim_thompson5910):

if you don't have a graphing calculator, you can use wolfram alpha http://www.wolframalpha.com/input/?i=s^3+-+7s^2+-+37+%3D+0